OpenStudy (michellemn):

Solve 4(2x - 1) - 3 = 9(x - 4). -29 32 13 29

OpenStudy (yanasidlinskiy):

Use Distributive Property first. \(\huge{8x-4-3=9x-32}\) Combine like terms \(\huge{8x-7=9x-32}\) Can you solve the rest on your own?

OpenStudy (anonymous):

x = 29

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

its easy

OpenStudy (yanasidlinskiy):

Direct answers are AGAINST the CoC. Please review the CoC again. Thank you.

OpenStudy (anonymous):

if you don't feel like going through all of that work all you have to do is substitue each number into the x values then you have to see if each side equals each other. if they do BINGO!!

OpenStudy (arabpride):

Alright - step by step :D 4(2x - 1) - 3 = 9(x - 4). Distributive Property: 4(2x - 1) - 3 = 9(x - 4). (4 * 2x) - (4 *1) - 3= (9 * x) - (9 *5) 8x - 4 - 3 = 9x - 4 Simplify: 8x - 4 - 3 = 9x - 4 (-4 -3) 8x -7 = 9x - 4 Combine Like terms: 8x -7 = 9x - 4 -8x -8x -7 = x - 4 +4 +4 x = -3

OpenStudy (arabpride):

Plug in -3 and you'll get your answer :D

OpenStudy (michellemn):

ok thanks guys

OpenStudy (arabpride):

You're Welcome *^-^

OpenStudy (michellemn):

@arabpride Can you help me with 1 more? If so, here Solve 4(x + 1) = 4x + 10. 0 6 All real numbers No solution

OpenStudy (arabpride):

All right same steps :D 4(x + 1) = 4x + 10 Distributive Property: 4(x + 1) = 4x + 10 (4 * x) + (4 * 1) = 4x + 10 4x + 4 = 4x + 10 Combine like terms 4x + 1 = 4x + 10 -1 -1 4x = 4x + 9 -4x -4x 0 ≠ 9 not sure - either 0 or no solution~

OpenStudy (michellemn):

ok thanks

OpenStudy (arabpride):

Welcome :D